# LDP for invariant measures of the Allen-Cahn equation with inhomogeneous boundary conditions and multiplicative noise

### v1

@@ 4.1 Exponential estimate
By the Sobolev embedding theorem, $W^{k^{\star},p^{\star}}(-L,L)$ is compactly embedded into $E$. We will prove the following proposition first.
@@
For any $r>0$, there exists $\epsilon_{r}>0$, $R_{r}>0$ such that
@@
$\|\bar{u}_{\epsilon}^{0}(t)\|_{W^{k^{\star},p^{\star}}}$ $=\|(-\Delta)^{\frac{k^{\star}}{2}}\bar{u}_{\epsilon}^{0}(t)\|_{L^{p^{\star}}}$
$\leq\int_{0}^{t}\left\|(-\Delta)^{\frac{k^{\star}}{2}}S(t-s)e^{-\lambda(t-s)}[F(\bar{u}_{\epsilon}^{0}(s))+\lambda\bar{u}_{\epsilon}^{0}(s)]\right\|_{L^{p^{\star}}}\,ds$
$\ \ \ \ +\epsilon^{\frac{1}{2}}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}$
$\leq\int_{0}^{t}(t-s)^{-\frac{k^{\star}}{2}}e^{-\lambda(t-s)}\;ds\cdot\sup_{s\in[0,t]}\|F(\bar{u}_{\epsilon}^{0}(s))+\lambda\bar{u}_{\epsilon}^{0}(s)\|_{L^{p^{\star}}}$
$\ \ \ \ +\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}.$
This implies that
$\begin{split}\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{W^{k^{\star},p^{\star}}}&\leq C\sup_{t\geq 0}\|F(\bar{u}_{\epsilon}^{0}(t))+\lambda\bar{u}_{\epsilon}^{0}(t)\|_{L^{p^{\star}}}+\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\\ &\leq C\sup_{t\geq 0}\left(\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{L^{3p^{\star}}}^{3}+\epsilon^{\frac{3}{2}}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{L^{3p^{\star}}}^{3}+1\right)\\ &\ \ \ \ +\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}.\end{split}$ (4.4)
@@
Next, we need to estimate $\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})\|_{L^{3p^{\star}}}^{3p^{\star}}$. Multiplying both parts of the equation (4.6) by $3p^{\star}\cdot Y_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}-1}(t)$ and integrating with respect to $\xi$ in [L,L][-L,L], since $p^{\star}>0$ is chosen to be an even number, by Young’s inequality, there exists $0<\sigma<1$ such that
@@
$=-3p^{\star}(3p^{\star}-1)\int_{-L}^{L}[D_{\xi}Y_{\lambda}(\bar{u}_{\epsilon}^{0})(t,\xi)]^{2}Y_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}-2}(t,\xi)\;d\xi$
@@
By Gronwall’s inequality, this implies that
$\sup_{t\geq 0}\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{L^{3p^{\star}}}^{3p^{\star}}\leq C\left(\sup_{t\geq 0}\epsilon^{\frac{3p^{\star}+2}{2}}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{L^{3p^{\star}+2}}^{3p^{\star}+2}+1\right).$ (4.7)
Thus, by (4.4), Sobolev inequality and (4.7),
$\mathbb{P}(\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r})$ $\leq\mathbb{P}\left(\sup_{t\geq 0}\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{L^{3p^{\star}}}^{3p^{\star}}\geq\left(\frac{R_{r}/3-C}{C}\right)^{p^{\star}}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{L^{3p^{\star}}}\geq\left(\frac{R_{r}}{3C}\right)^{\frac{1}{3}}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\frac{R_{r}}{3}\right)$
$\leq\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\left[\frac{\left(\frac{R_{r}/3-C}{C}\right)^{p^{\star}}-C}{C}\right]^{\frac{1}{3p^{\star}+2}}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\left(\frac{R_{r}}{3C}\right)^{\frac{1}{3}}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\frac{R_{r}}{3}\right)$
$\leq 3\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r}\right).$
Here, $R_{r}$ is a universal constant that may vary from line to line but depends only on $r$. Hence,
@@
$\leq\mathbb{P}(\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r})$
$\leq 3\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r}\right).$
@@
For any $r>0$, there exist $\epsilon_{r}>0$ and $R_{r}>0$ such that
$\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r}\right)\leq\exp(-\frac{r}{\epsilon}),\;\epsilon<\epsilon_{r}.$ (4.8)
@@
$\ \ \ \ \mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r}\right)$
$=\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r},\;\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\leq\bar{R}_{r}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r},\;\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}>\bar{R}_{r}\right)$
$\leq\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r},\;\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\leq\bar{R}_{r}\right)$
$\ \ \ \ +\mathbb{P}\left(\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\geq\bar{R}_{r}\right).$
By the factorization formula,
@@
$\begin{split}\|\gamma_{\lambda}(v)(t)\|_{W^{k^{\star},p^{\star}}}^{p^{\star}}&\leq C_{\alpha}^{p^{\star}}\left(\int_{0}^{t}(t-s)^{\alpha-1-\frac{k^{\star}}{2}}e^{-\lambda(t-s)}\|\Gamma^{\alpha}(v)(s)\|_{L^{p^{\star}}}\;ds\right)^{p^{\star}}\\ &\leq C_{\alpha}^{p^{\star}}\left(\int_{0}^{t}(t-s)^{(\alpha-1-\frac{k^{\star}}{2})\frac{p^{\star}}{p^{\star}-1}}e^{-\frac{\lambda}{2}(t-s)\frac{p^{\star}}{p^{\star}-1}}\;ds\right)^{(p^{\star}-1)}\\ &\ \ \ \ \times\left(\int_{0}^{t}e^{-\frac{\lambda}{2}(t-s)p^{\star}}\|\Gamma^{\alpha}(v)(s)\|_{L^{p^{\star}}}^{p^{\star}}\;ds\right)\\ &\leq C\times\left(\int_{0}^{t}e^{-\frac{\lambda}{2}(t-s)p^{\star}}\;ds\right)\times\sup_{s\leq t}\|\Gamma^{\alpha}(v)(s)\|_{L^{p^{\star}}}^{p^{\star}}\\ &\leq C\sup_{s\leq t}\|\Gamma^{\alpha}(v)(s)\|_{L^{p^{\star}}}^{p^{\star}}.\end{split}$ (4.9)
Thus, for any $\|v\|_{C([0,+\infty);E)}\leq\bar{R}_{r}$,
$\sup_{t\geq 0}\|\gamma_{\lambda}(v)(t)\|_{W^{k^{\star},p^{\star}}}\leq C\sup_{t\geq 0}\left\|\Gamma^{\alpha}_{\bar{R}_{r}}(t)\right\|_{L^{p^{\star}}},$ (4.10)
@@
$\mathbb{E}\left[\sup_{s\leq t}\|\Gamma^{\alpha}_{\bar{R}_{r}}(s)\|_{L^{p^{\star}}}^{p^{\star}}\right]$ $\leq\mathbb{E}\int_{-L}^{L}\sup_{s\leq t}|\Gamma^{\alpha}_{\bar{R}_{r}}(s,\xi)|^{p^{\star}}\;d\xi$
$\leq\int_{-L}^{L}(\mathbb{E}|\Gamma^{\alpha}_{\bar{R}_{r}}(t,\xi)|^{2})^{\frac{p^{\star}}{2}}\;d\xi$
$\leq C_{\alpha}^{\frac{p^{\star}}{2}}(1+\bar{R}_{r})^{p^{\star}}.$
Taking $t\rightarrow\infty$ on the left hand side of the above inequality and by monotone convergence, we obtain that
$\mathbb{E}\left[\sup_{0\leq t<\infty}\|\Gamma^{\alpha}_{\bar{R}_{r}}(t)\|_{L^{p^{\star}}}^{p^{\star}}\right]\leq C_{\alpha}^{\frac{p^{\star}}{2}}(1+\bar{R}_{r})^{p^{\star}}.$
@@
$\leq\frac{\epsilon^{\frac{p^{\star}}{2}}C_{\alpha}^{p^{\star}}(1+\bar{R}_{r})^{p^{\star}}}{\delta^{p^{\star}}}.$
Then, for any $\delta>0$
$\lim_{\epsilon\rightarrow 0}\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\Gamma^{\alpha}_{\bar{R}_{r}}(t)\|_{L^{p^{\star}}}\geq\delta\right)=0.$ (4.12)
By Theorem 2.10 and Theorem 2.13 of [salins2019equivalences], (4.12) implies that the family of measures
$\upsilon^{\epsilon}(B):=\mathbb{P}\left(\epsilon^{\frac{1}{2}}\Gamma^{\alpha}_{\bar{R}_{r}}(\cdot)\in B\right),\;B\in\mathcal{B}\left(C([0,+\infty);L^{p^{\star}})\right),$
satisfies LDP in $L^{p^{\star}}$ with rate functional
$I^{\alpha}_{\bar{R}_{r}}(z)=\frac{1}{2}\inf\left\{\int_{0}^{+\infty}\|f(t)\|_{H}^{2}\;dt;z=\Gamma_{\bar{R}_{r}}^{\alpha}(f)\right\},$
@@
$\Gamma_{\bar{R}_{r}}^{\alpha}(f)(t):=\int_{0}^{t}(t-s)^{-\alpha}S(t-s)e^{-\lambda(t-s)}(1+\bar{R}_{r})f(s)\;ds.$
@@
$K^{\alpha}_{\bar{R}_{r}}(r):=\{z\in C([0,+\infty);E):I^{\alpha}_{\bar{R}_{r}}(z)\leq r\}$
to be level sets of $I^{\alpha}_{\bar{R}_{r}}$. By the same procedure as above, for $f\in L^{2}([0,+\infty);H)$ and $t>0$, $\|\Gamma_{\bar{R}_{r}}^{\alpha}(f)\|_{C([0,+\infty);L^{p^{\star}})}\leq C\|f\|_{L^{2}([0,+\infty);H)}$. This implies that for any $r>0$, $K^{\alpha}_{\bar{R}_{r}}(r)$ is bounded in $L^{p^{\star}}$. Then, by the same argument as Theorem 3.9 in [BAI2026111284] and (4.10), there exists $\delta^{\star}>0$ and $\epsilon_{r}$, $R_{r}$ depends on $\bar{R}_{r}$, $r$, such that
$\ \ \ \ \mathbb{P}\left(\bigcup_{\|v\|_{C([0,+\infty);E)}\leq\bar{R}_{r}}\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(v)(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r}\right)$
$\leq\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\Gamma^{\alpha}_{\bar{R}_{r}}(t)\|_{L^{p^{\star}}}\geq R_{r}\right)$
$\leq\mathbb{P}\left(\epsilon^{\frac{1}{2}}\Gamma^{\alpha}_{\bar{R}_{r}}(\cdot)\in B_{C([0,+\infty);L^{p^{\star}})}^{c}(K^{\alpha}_{\bar{R}_{r}}(r),\delta)\right)$
@@
Hence, for $\epsilon\leq\epsilon_{r}$,
$\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r},\;\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\leq\bar{R}_{r}\right)\leq\exp(-\frac{r}{\epsilon}).$ (4.13)

### v2

@@ 4.1 Exponential estimates
By the Sobolev embedding theorem, $W^{k^{\star},p^{\star}}(-L,L)$ is compactly embedded into $E$. We will first prove the following proposition.
@@
For any $r>0$, there exist $\epsilon_{r}>0$ and $R_{r}>0$ such that
@@
$\begin{split}&\ \ \ \ \|\bar{u}_{\epsilon}^{0}(t)\|_{W^{k^{\star},p^{\star}}}=\|(-\Delta)^{\frac{k^{\star}}{2}}\bar{u}_{\epsilon}^{0}(t)\|_{L^{p^{\star}}}\\ &\leq\int_{0}^{t}\left\|(-\Delta)^{\frac{k^{\star}}{2}}S(t-s)e^{-\lambda(t-s)}[F(\bar{u}_{\epsilon}^{0}(s))+\lambda\bar{u}_{\epsilon}^{0}(s)]\right\|_{L^{p^{\star}}}\,ds\\ &\ \ \ \ +\epsilon^{\frac{1}{2}}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\\ &\leq\int_{0}^{t}(t-s)^{-\frac{k^{\star}}{2}}e^{-\lambda(t-s)}\;ds\cdot\sup_{s\in[0,t]}\|F(\bar{u}_{\epsilon}^{0}(s))+\lambda\bar{u}_{\epsilon}^{0}(s)\|_{L^{p^{\star}}}\\ &\ \ \ \ +\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\\ &\leq C\sup_{s\in[0,t]}\|F(\bar{u}_{\epsilon}^{0}(s))+\lambda\bar{u}_{\epsilon}^{0}(s)\|_{L^{p^{\star}}}+\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\\ &\leq C\sup_{s\in[0,t]}\left(\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{3p^{\star}}}^{3}+\epsilon^{\frac{3}{2}}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{3p^{\star}}}^{3}+\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}+1\right).\end{split}$ (4.4)
@@
Next, we need to estimate $\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})\|_{L^{3p^{\star}}}^{3p^{\star}}$. Multiplying both sides of equation (4.6) by $3p^{\star}Y_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}-1}(t)$ and integrating with respect to $\xi$ in [L,L][-L,L], since $p^{\star}>0$ is chosen to be an even number, by Young’s inequality, there exists $0<\sigma<1$ such that
@@
$=-3p^{\star}(3p^{\star}-1)\int_{-L}^{L}[\nabla Y_{\lambda}(\bar{u}_{\epsilon}^{0})(t,\xi)]^{2}Y_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}-2}(t,\xi)\;d\xi$
@@
By Grönwall’s inequality, this implies that
$\sup_{s\in[0,t]}\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{3p^{\star}}}^{3p^{\star}}\leq C\left(\sup_{s\in[0,t]}\epsilon^{\frac{3p^{\star}+2}{2}}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{3p^{\star}+2}}^{3p^{\star}+2}+1\right).$ (4.7)
Thus, by (4.4), the Sobolev inequality and (4.7),
$\mathbb{P}(\|\bar{u}_{\epsilon}^{0}(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r})$ $\leq\mathbb{P}\left(\sup_{s\in[0,t]}\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{3p^{\star}}}^{3p^{\star}}\geq\left(\frac{R_{r}/3-C}{C}\right)^{p^{\star}}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{3p^{\star}}}\geq\left(\frac{R_{r}}{3C}\right)^{\frac{1}{3}}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\frac{R_{r}}{3}\right)$
$\leq\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\left[\frac{\left(\frac{R_{r}/3-C}{C}\right)^{p^{\star}}-C}{C}\right]^{\frac{1}{3p^{\star}+2}}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\left(\frac{R_{r}}{3C}\right)^{\frac{1}{3}}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\frac{R_{r}}{3}\right)$
$\leq 3\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r}\right).$
Here, $\tilde{R}_{r}$ is a constant derived from $R_{r}$. Hence,
@@
$\leq\lim_{T\rightarrow\infty}\frac{3}{T}\int_{0}^{T}\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r}\right)\;dt.$
@@
For any $r>0$, there exist $\epsilon_{r}>0$ and $\tilde{R}_{r}>0$ such that
$\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r}\right)\;dt\leq\exp(-\frac{r}{\epsilon}),\;\epsilon<\epsilon_{r}.$ (4.8)
@@
$\ \ \ \ \mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r}\right)$
$=\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r},\;\sup_{s\in[0,t]}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\leq\bar{R}_{r}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r},\;\sup_{s\in[0,t]}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}>\bar{R}_{r}\right)$
$\leq\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r},\;\sup_{s\in[0,t]}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\leq\bar{R}_{r}\right)$
$\ \ \ \ +\mathbb{P}\left(\sup_{s\in[0,t]}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\geq\bar{R}_{r}\right).$
By the factorisation formula,
@@
$\begin{split}\|\gamma_{\lambda}(v)(t)\|_{W^{k^{\star},p^{\star}}}^{p^{\star}}&\leq C_{\alpha}^{p^{\star}}\left(\int_{0}^{t}(t-s)^{\alpha-1-\frac{k^{\star}}{2}}e^{-\lambda(t-s)}\|\Gamma^{\alpha}(v)(s)\|_{L^{p^{\star}}}\;ds\right)^{p^{\star}}\\ &\leq C_{\alpha}^{p^{\star}}\left(\int_{0}^{t}(t-s)^{(\alpha-1-\frac{k^{\star}}{2})\frac{p^{\star}}{p^{\star}-1}}e^{-\frac{\lambda}{2}(t-s)\frac{p^{\star}}{p^{\star}-1}}\;ds\right)^{(p^{\star}-1)}\\ &\ \ \ \ \times\left(\int_{0}^{t}e^{-\frac{\lambda}{2}(t-s)p^{\star}}\|\Gamma^{\alpha}(v)(s)\|_{L^{p^{\star}}}^{p^{\star}}\;ds\right)\\ &\leq C\times\left(\int_{0}^{t}e^{-\frac{\lambda}{2}(t-s)p^{\star}}\;ds\right)\times\sup_{s\in[0,t]}\|\Gamma^{\alpha}(v)(s)\|_{L^{p^{\star}}}^{p^{\star}}\\ &\leq C\sup_{s\in[0,t]}\|\Gamma^{\alpha}(v)(s)\|_{L^{p^{\star}}}^{p^{\star}}.\end{split}$ (4.9)
Thus, for any $\|v\|_{C([0,t];E)}\leq\bar{R}_{r}$,
$\sup_{s\in[0,t]}\|\gamma_{\lambda}(v)(s)\|_{W^{k^{\star},p^{\star}}}\leq C\sup_{s\in[0,t]}\left\|\Gamma^{\alpha}_{\bar{R}_{r}}(t)\right\|_{L^{p^{\star}}},$ (4.10)
@@
$\mathbb{E}\left[\sup_{s\in[0,t]}\|\Gamma^{\alpha}_{\bar{R}_{r}}(s)\|_{L^{p^{\star}}}^{p^{\star}}\right]$ $\leq\mathbb{E}\int_{-L}^{L}\sup_{s\leq t}|\Gamma^{\alpha}_{\bar{R}_{r}}(s,\xi)|^{p^{\star}}\;d\xi$
$\leq\int_{-L}^{L}C_{p^{\star}}(\mathbb{E}|\Gamma^{\alpha}_{\bar{R}_{r}}(t,\xi)|^{2})^{\frac{p^{\star}}{2}}\;d\xi$
$\leq C_{p^{\star}}C_{\alpha}^{\frac{p^{\star}}{2}}(1+\bar{R}_{r})^{p^{\star}}.$
Taking $t\rightarrow\infty$ on the left hand side of the above inequality and by the monotone convergence theorem, we obtain that
$\mathbb{E}\left[\sup_{t\geq 0}\|\Gamma^{\alpha}_{\bar{R}_{r}}(t)\|_{L^{p^{\star}}}^{p^{\star}}\right]\leq C_{\alpha,p^{\star}}(1+\bar{R}_{r})^{p^{\star}}.$
@@
$\leq\frac{\epsilon^{\frac{p^{\star}}{2}}C_{\alpha,p^{\star}}(1+\bar{R}_{r})^{p^{\star}}}{\delta^{p^{\star}}}.$
Then, taking any $\delta,\eta>0$ and $t\geq 0$ there exists $\epsilon_{r}>0$ such that for any $\epsilon<\epsilon_{r}$,
$\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\Gamma^{\alpha}_{\bar{R}_{r}}(t)\|_{L^{p^{\star}}}\geq\delta\right)<\eta.$ (4.12)
Here, $\epsilon_{r}$ does not depend on $t$. By Theorem 2.10 and Theorem 2.13 of [salins2019equivalences], (4.12) implies that the family of measures
$\upsilon^{\epsilon}(B):=\mathbb{P}\left(\epsilon^{\frac{1}{2}}\Gamma^{\bar{R}_{r},\alpha}_{t}(\cdot)\in B\right),\;B\in\mathcal{B}\left(C([0,t];L^{p^{\star}})\right),$
satisfies LDP in $L^{p^{\star}}$ uniformly for $t\geq 0$ with the rate functional
$I^{\bar{R}_{r},\alpha}_{t}(z)=\frac{1}{2}\inf\left\{\int_{0}^{t}\|f(s)\|_{H}^{2}\;dt;z=\Gamma^{\bar{R}_{r},\alpha}_{t}(f)\right\},$
@@
$\Gamma^{\bar{R}_{r},\alpha}_{t}(f)(\tau):=\int_{0}^{\tau}(\tau-s)^{-\alpha}S(\tau-s)e^{-\lambda(\tau-s)}(1+\bar{R}_{r})f(s)\;ds.$
@@
$K^{\bar{R}_{r},\alpha}_{t}(r):=\{z\in C([0,t];E):I^{\bar{R}_{r},\alpha}_{t}(z)\leq r\}$
to be a level set of $I^{\bar{R}_{r},\alpha}_{t}$. By the same procedure as above, for $f\in L^{2}(0,t;H)$ and $t\geq 0$, $\|\Gamma^{\bar{R}_{r},\alpha}_{t}(f)\|_{C([0,t];L^{p^{\star}})}\leq C\|f\|_{L^{2}(0,t;H)}$. This implies that for any $r>0$, $K^{\bar{R}_{r},\alpha}_{t}(r)$ is bounded in $C([0,t];L^{p^{\star}})$. Then, by the same argument as Theorem 3.9 in [BAI2026111284] and (4.10), there exist $\delta^{\star}>0$ and $\epsilon_{r}$, $\tilde{R}_{r}$ depend on $\bar{R}_{r}$, $r$ such that for any $t\geq 0$
$\ \ \ \ \mathbb{P}\left(\bigcup_{\|v\|_{C([0,t];E)}\leq\bar{R}_{r}}\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(v)(s)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r}\right)$
$\leq\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\Gamma^{\bar{R}_{r},\alpha}_{t}(s)\|_{L^{p^{\star}}}\geq\tilde{R}_{r}\right)$
$\leq\mathbb{P}\left(\epsilon^{\frac{1}{2}}\Gamma^{\bar{R}_{r},\alpha}_{t}(\cdot)\in B_{C([0,t];L^{p^{\star}})}^{c}(K^{\bar{R}_{r},\alpha}_{t}(r),\delta)\right)$
@@
Hence, for $\epsilon\leq\epsilon_{r}$ and all $t\geq 0$
$\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r},\;\sup_{s\in[0,t]}\|\bar{u}_{\epsilon}^{0}(s)\|_{E}\leq\bar{R}_{r}\right)\leq\exp(-\frac{r}{\epsilon}).$ (4.13)